OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with an x-intercept of 6 and a y-intercept of 5.

OpenStudy (savvy):

\[y = -5x/6 + 5\]

OpenStudy (anonymous):

y = -6x + 5 ?

OpenStudy (savvy):

no the eqn. you wrote can be written in intercept form as (y/5) + (x/(-5/6)) = 1 which means line has an x int. = -5/6 which is against the question....

OpenStudy (anonymous):

y = -5/6 x + 5...

OpenStudy (savvy):

yep.....

OpenStudy (anonymous):

Thank you (: